a 1200kg car coasts from rest down a driveway that is inclined 20degrees. to the horizontal and is 15m long. how fast is the car going at the end of the delivery if....
(a) friction is negligible
(book answer is 10m/s)

(b) a friction force of 3000n opposes the motion?
(book answer is 5.1 m/s)


dont know where to begin to solve these 2questions. I tried differant formulas, but nothing I tried came out to the answers in the book.
help
Answer #1
Many if not most problems in physics are not solved by plugging directly into a single equation. It is better to prceed in logical steps using formulas you are familiar with.

In this case, the place to start is with conservation of energy. The vertical drop of the runway is
H = 15 sin 20 = 5.13
The Potential energy loss MgH equals the Kinetic Energy gain, so
MgH = (1/2) M V^2
V = sqrt(2 g H) = 100.3 m/s
(I used 9.81 m/s^2 for g)
Note that you did not need to use the mass for this one.

To do the second case, subtract friction work done
3000N*15 m = 45,000 J from the potential energy decrease, and solve for V again.

M g H - 45,000 = (1/2) M V^2
You will need to use the mass this time.
answered by: Palesa
Answer #2
see this very related problem:
http://www.jiskha.com/display.cgi?id=1266960070
answered by: Jaer
Answer #3
Two long, parallel wires separated by 30 cm each carry currents of 4.0 A in a horizontal direction.

B)Find if they are in opposite directions.
answered by: DELLI